two out of 6 or one third
Probability of rolling a 3 is 2/36 & the probability of rolling a 6 is 5/36; probability of rolling a 3 or 5 is 2/36 + 5/36 = 7/36.
a 2 out of 6 probability
The probability of not rolling a sum of six with two fair dice is 1 minus the probability of rolling a sum of six. There are 36 permutations of rolling two dice. Of these, five sum to six, 1+5, 2+4, 3+3, 4+2, and 5+1. The probability, then of rolling a sum of six is 5 in 36. The probability, then of not rolling a sum of six is 31 in 36, or about 0.8611.
5 out of 6 or 83.333%
The probability of rolling a sum of 5 with two die is 4 over 36, or 2 over 18, or 1 over 9.
5/36, or 14 percent.
The probability is 11/36. There are 36 possible outcomes, but 25 of these don't involve a 2 on either die. The way to answer this is to ask what is the probability of not rolling a 2 on either die and subtract this from one. The probability of not rolling a 2 with the first die is 5/6. It is of course also 5/6 chance of not rolling a 2 with the second die. So we multiply the probabilities and subtract from 1. 1 - (5/6 * 5/6) = 1 - (25/36) = 11/36.
The probability is 1/6 x 5=5/12
Firstly, the probability when tossing a coin and getting a head or tail is 1/2, then rolling a die, there are 6 sides so the chance of rolling any number is 1/6, there are 2 chances of rolling greater than 4 ie 5 and 6, so the probability of rolling a 5 or 6 in 1/3, as these are independent events you multiply the probability getting a heads of tails, (1/2) by the probability of rolling a five or six, (1/3) which gives you 1/6 or 0.1666 recurring.
The probability of a head is 1/2. The probability of an odd number is 1/2. 1/2 x 1/2 = 1/4 or 0.25
5/216
If you rolled 2 fair dice, the probability of having a sum of 6 is 5 over 36
Two thirds (2/3)
There could be many questions: What is the probability of rolling an even number. What is the probability of rolling an odd number. What is the probability of rolling a number less than 4. What is the probability of rolling a number more than 3. What is the probability of rolling 1,4, or 6. Basically it could be any question about the probability of rolling half of the faces.
.33333333333333333
1/2
1/2.